Five percent (0.05) of the parts produced by a machine are defective. Twenty(20) parts are selected at random for study.

A. what is the probability that more than 4 parts are defective?

B. what is the probability that exactly 3 parts are defective?

C. what is the expected number of defective parts in the sample?

D. What is the variance and standard deviation of defective parts in the sample?
Answer #1

Solution Giren p=oos n=20 XBCnP) P(X=x)=Q) p (p) nok ; x=0, 1,2 p(X24) = 1-pcxct) =)-4p XEO) +PCX=1) + P(x=2) + PCX= 3) + p(xE(X)= 4= np = 20x0.05 = ) | E(X)=1 I is the expected number of defective parts in the sample. varix) = n p q 0 - 1 9= )-P = 2